# From Actions to Words: Towards Abstractive-Textual Policy Summarization in RL

### v2

@@ 6 Expert Evaluation of Summaries
4 0.750 0.731 0.914 0.795
5 0.917 0.808
6 1.000 0.846
Crafter
Resource-Collector
7 0.938 0.893 0.931 0.871
8 0.938 0.857
9 0.917 0.864
Crafter
Random
7 0.929 0.917 0.902 0.929
8 1.000 0.958
9 0.778 0.857
Table 5: Recall and Precision of Expert Insights for SySLLM.
7 User Study
@@
To evaluate the usefulness of the SySLLM summaries, we conducted a user study comparing SySLLM summaries to HIGHLIGHTS-DIV (hereafter referred to as “HIGHLIGHTS”) summaries, which is a standard video-based benchmark in XRL [17]. For the study, we focused on the three agents from the MiniGrid Unlock environment: goal-directed agent, short-sighted agent, and random agent. These agents were chosen to represent qualitatively distinct behaviors, ensuring a diverse set of policies for participants to evaluate. This variety allowed us to better assess how well the summaries captured different decision-making strategies and provided interpretable insights into the agents’ behavior. For completeness, we provide an intuitive explanation of HIGHLIGHTS in A. We assessed participants’ subjective preferences as well as their ability to identify agent behavior based on the summaries.
7.1 Procedure
Participants were first shown a tutorial explaining the rules of the Unlock environment. They then had to pass a quiz ensuring they read and understood the rules. Next, they were asked to complete two different tasks.
Task 1: Summary Preferences
In the first task, participants evaluated and compared summaries of agents based on their policies. Initially, they watched a 120-second video showcasing full episodes to familiarize themselves with the agent’s policy. Following this, participants were presented with the first summary, either from the SySLLM or HIGHLIGHTS condition (see Section 7.2), and rated their agreement on a series of metrics using a 7-point Likert scale. The metrics, adapted from Hoffman’s explanation satisfaction questionnaire [55], included Clarity, Understandability, Completeness, Satisfaction, Usefulness, Accuracy, Improvement, and Preference. These adaptations were tailored to the study’s specific needs and excluded irrelevant items (see G for the full scale).
Participants then evaluated the second summary, corresponding to the alternative condition (HIGHLIGHTS or SySLLM), using the same set of questions. This allowed for a direct comparison between the two summary types. After assessing both summaries, participants answered additional questions that compared the summaries and expressed their preferences, highlighting which summary they found more helpful and explaining their reasons.
Task 2: Agent Identification
While the first task evaluated participants’ subjective perceptions of the summaries, the second task assessed their objective ability to match a given summary to the correct agent. Participants were presented with summaries, each linked to either the SySLLM or HIGHLIGHTS condition, and shown three short experience videos (approximately 20 seconds each) depicting different agents’ actions in the environment. They were asked to identify which video corresponded to the behavior described in the summary. Participants indicated their chosen video and rated their confidence on a 1-7 Likert scale, along with justifications for their decision in an open text format.
To keep the experiment concise and focused on key comparisons, we set up three matching questions—labeled $Q1$, $Q2$, and $Q3$. Each question featured a summary of a different agent: the goal-directed agent in $Q1$, the random agent in $Q2$, and the short-sighted agent in $Q3$. Depending on the condition to which a participant was assigned, they encountered either a text-based summary (SySLLM) or a video summary (HIGHLIGHTS) for these agents. In each question, alongside the video of the associated agent, participants were also shown two additional videos featuring different, unrelated agents. The task for the participants was to correctly identify the agent that matched the given summary. Success was measured by the participant’s ability to accurately associate the correct agent with its respective summary.
7.2 Experimental Conditions
Participants were divided into four conditions, varying by summary type (SySLLM or HIGHLIGHTS) and agent type (goal-directed or short-sighted). For Task 1, we used a within-subject design where participants viewed summaries of an agent’s behavior in two sequences. To counter ordering effects, one group received the SySLLM summary first and the HIGHLIGHTS summary second, while the other group received the reverse order. In Task 2, we employed a between-subject design. Participants viewed only SySLLM or HIGHLIGHTS summaries, corresponding to the sequence they were exposed to in Task 1. This setup ensured that participants experienced one of four unique combinations of summary and agent types in a controlled manner. Details of these conditions are shown in Table 6.
Benchmark
For generating the HIGHLIGHTS summaries, we used the following parameters: 300 traces (episodes), a context length of 5, and a total of 20 highlights to effectively capture key behaviors and the necessary context within the Unlock environment. These choices ensure that the summaries provide a detailed yet concise depiction of critical actions, appropriately matching the environment’s scale and complexity.
@@ 4 HIGHLIGHTS $\rightarrow$ SySLLM Short-sighted Agent HIGHLIGHTS
Table 6: Experimental conditions.
7.3 Participants
We recruited participants through Prolific (N = 200). Participants were native English speakers from the US, UK, Australia, or Canada. Participants received a 3.75 pound base payment and an additional, 1 pound bonus if participants answered all questions correctly in Task 2. Participants who failed the attention question were excluded from the final analysis, as well as participants who completed the survey in less time than the length of the videos presented. After exclusions, we had 192 participants (94 female, Mean age = 36.41, STD = 12.05).
7.4 Results
Task 1: Summary Preferences
The SySLLM summaries consistently received higher ratings compared to the HIGHLIGHTS summaries across all metrics (see Figure 4). To assess the statistical significance of the differences in ratings, we first calculate the average scores for each participant under both the SySLLM and HIGHLIGHTS conditions. We then compute the difference between these averages for each participant. A paired t-test is conducted on these differences to evaluate whether the SySLLM condition significantly outperforms the HIGHLIGHTS condition across all metrics. The results indicate a statistically significant difference in favor of the SySLLM condition, with statistic $T=13.99$ and $p<10^{-33}$.
Following the evaluations of specific summaries, we asked the participants to subjectively compare the textual and video summaries. Specifically, we asked them which summary they think is better for understanding the agent’s overall behavior based on the observed video of a large sample of trajectories. They were further asked to rate the extent to which they preferred the textual (or visual) summary (on a 1-7 Likert scale). The results we obtained are conclusive. The responses indicate a clear preference for the SySLLM summary, with 75.5% of participants favoring the language-based summaries. Regarding the second question, the mean rating was 5.97 with a standard deviation of 1.44. In the 1-7 Likert scale, a score of 4 represents indifference to the different policies; hence, a mean score of 5.97 indicates a substantial preference for the textual summary.
Qualitative Analysis The open text feedback we received from participants of Task 1 turned out to be highly valuable. Some participants explained that while HIGHLIGHTS summaries required their own interpretation of the videos, SySLLM summaries provided an interpretation of the agent’s policy, e.g., “the [SySLLM] summary was thorough in explaining the movement of the agent. It describes information that isn’t visible in the video. The information in the summary was objective while watching the video summary and interpreting it is based on my view which is subjective.” Other participants mentioned that SySLLM summaries also explained why the agent acted as it did. For instance, one participant noted that “there are instances in the video where the agent seems to turn random corners instead of straight as you would expect it to. The summary explains why.” Another participant wrote that “the video itself does not fully explain but rather how you interpret what it is doing and in the case of errors it starts doing something wrong you don’t know what causes it”.
Additionally, some participants liked that SySLLM summaries provided more context, e.g., “the textual summary provides context for what is occurring. Conversely, the video summary doesn’t inform the viewer of the ’behind the scenes’ that is occurring. From the textual information, I identified many strategies and code that the agent is using such as using the maze strategy to always turn right when it hits a wall. On the other hand, the video does not provide this information unless you have prior knowledge such as coding experience on conditionals and iterations to know what it occurring.”
@@ 7.4 Results
Figure 4: Participant scores for Task 1. Each metric includes a box plot of the Likert scores. Notably, SySLLM receives higher ratings than HIGHLIGHTS across all metrics.
Task 2: Agent Identification
As shown in Figure 5, participants’ ability to correctly identify the agent based on the provided summaries varied between the SySLLM and HIGHLIGHTS conditions across the three questions.
For $Q1$ ($\chi^{2}=1.0523$, $p=0.3050$, odds ratio (OR) = 0.5843), $Q2$ ($\chi^{2}=0.6703$, $p=0.4129$, OR = 0.7550), and $Q3$ ($\chi^{2}=0.7279$, $p=0.3936$, OR = 1.3875), the Chi-Square tests showed no statistically significant differences. Independent t-tests showed no statistically significant differences in confidence scores between the SySLLM and HIGHLIGHTS conditions for $Q1$ (t=0.6202t=-0.6202, $p=0.5359$) and $Q2$ ($t=0.5565$, $p=0.5785$). However, for $Q3$, there was a statistically significant difference ($t=3.4197$, $p=0.0008$).
To conclude, participants did not show a significant preference for the SySLLM or HIGHLIGHTS condition when identifying the correct agent based on the summaries provided, except for $Q3$ where the SySLLM condition had significantly higher confidence scores. Further investigation may be required to understand the nuances behind these results and whether different summary methods or conditions might yield different outcomes.
Figure 5: Correctness rate in Task 2. For each question, we measure the proportion of users that identified the correct agent under both SySLLM, HIGHLIGHTS, and a random guess. For each bar, we included the 95% confidence intervals. Notice that both SySLLM and HIGHLIGHTS receive higher scores than the random guess. For all three questions, the difference in correctness rate between SySLLM and HIGHLIGHTS is not statistically significant.
8 Discussion and Future Work
In this paper, we introduced SySLLM, an innovative framework leveraging large language models to synthesize reinforcement learning policies into clear, structured narratives. Our method addresses key shortcomings of traditional extractive-visual methods by moving away from fragmented, scenario-specific snapshots toward holistic, language-based summaries. This shift significantly reduces the effort required for users to interpret complex agent behaviors, facilitating better comprehension of strategic decisions and adaptability. Our evaluations provide strong evidence that SySLLM effectively captures nuanced policy behaviors and aligns closely with expert-generated summaries. The qualitative alignment with human experts highlights the capability of language models to accurately convey intricate behavioral patterns and strategic nuances, reinforcing the effectiveness of a textual-abstractive approach. Furthermore, user preference studies demonstrate clear advantages in interpretability and user engagement over visual-based summarization methods.
The broader implications of SySLLM suggest promising new directions for research and practice. We advocate for increased exploration into abstractive-textual explanations in global explanation in reinforcement learning, encouraging the community to investigate advanced techniques such as fine-tuning, in-context learning, and integration with reasoning frameworks. Embracing this paradigm can lead to significant advancements in how we communicate complex policies, ultimately enhancing trust, transparency, and adoption of reinforcement learning in critical real-world scenarios.
Captioners
Our approach relies on the availability of effective captioning functions, a common requirement in language-integrated methodologies as seen in prior research [35, 46, 45]. In simulated environments, this dependency is relatively straightforward to address, as ground truth states are accessible, and many RL environments—such as NetHack/MiniHack [56, 57] and text-based games [58, 59]—already incorporate language-based features. For more complex or real-world scenarios, however, implementations would require robust object detection [60], advanced captioning frameworks [61], action recognition systems [62], or vision-language models [63]. Alternatively, annotator models [35] or pretrained foundation models [64] could generate captions. The continued advancement of general-purpose captioning models suggests that off-the-shelf solutions will soon support an even broader range of tasks, enhancing the practicality of our approach.
LLMs Limitations
The current performance of SySLLM is bounded by the capabilities and limitations of the underlying language models. Context window constraints restrict the amount of agent trajectory data that can be analyzed at once, potentially limiting the ability to identify long-term patterns or rare behaviors. Additionally, LLMs can occasionally generate plausible-sounding but incorrect explanations (hallucinations), particularly when faced with unusual or complex agent behaviors. As LLM technology evolves, particularly with expanded context windows and improved factuality, we anticipate that these limitations will diminish, enabling more comprehensive and reliable agent behavior summaries.
Building on SySLLM’s strengths while addressing its limitations, we identify several promising directions for future research:
Interactive Question-Answering Framework
A promising avenue for future work is extending SySLLM to support an interactive question-answering (QA) framework. This would enable users to directly query the agent’s behavior, asking targeted questions such as “Why did the agent choose action A over B in this situation?” or “How would the agent respond to condition X?” Such interactivity could transform SySLLM from a static summarization tool into a dynamic interface for policy exploration, offering a more flexible and intuitive means of understanding agent behavior. Prior work on logical querying of agents, such as [65], highlights the feasibility and potential of QA-based explanations in RL. Inspired by these developments, we envision a natural language-based QA interface grounded in narrative trajectories, allowing users to probe specific aspects of agent decision-making in a more accessible way.
Multi-modal Summaries
Integrating natural language explanations with visual representations would significantly enhance the communicative power of SySLLM. For each behavioral pattern identified by the language model, the system could provide corresponding visual evidence extracted from agent traces such as state transitions, action sequences, or critical decision points. This multi-modal approach would ground abstract explanations in concrete examples, potentially mitigating hallucination issues while making complex policies more accessible to users with diverse backgrounds and expertise levels.
Comparative Analysis Framework
Future work could extend SySLLM to automatically compare multiple agents’ behaviors, identifying similarities, differences, and relative strengths. This comparative capability would be particularly valuable for evaluating alternative training approaches, understanding the effects of hyperparameter changes, or analyzing how agent behavior evolves during training. By articulating these comparisons in natural language, SySLLM could provide intuitive insights into the factors that influence agent performance and behavior.
Acknowledgments
This project was funded by the ERC Starting Grant #101078158 CONVEY and the Israel Science Foundation (ISF) under Grant No. 3079/24.

### v3

@@ 6. Illustrative Policy Summaries
  1. (1)
    The agent moves towards the green goal once it enters its field of vision, adjusting its path accordingly.
  2. (2)
    It avoids collisions with walls through timely directional changes.
Dynamic Obstacles
  1. (1)
    The agent effectively avoids moving obstacles (blue balls) by adjusting its movement.
  2. (2)
    It identifies objects in its field of vision and makes informed navigation decisions.
  3. (3)
    In dense obstacle scenarios, occasional inefficiencies or unnecessary turns are observed.
Crafter – Resource-collector
  1. (1)
    Strong Focus on Resource Collection: The agent consistently prioritizes gathering essential resources such as wood, stone, and drink, which are foundational for crafting tools and maintaining basic survival metrics.
  2. (2)
    Effective Basic Tool Crafting: Regular crafting of basic tools like wood pickaxes and swords enables the agent to enhance resource collection and engage in occasional combat.
  3. (3)
    Achievement Unlocking: The agent reliably unlocks achievements related to resource collection and basic tool crafting but struggles with more advanced achievements, highlighting a potential area for improvement.
  4. (4)
    Moderate Combat Engagement: The agent occasionally engages with zombies, using crafted tools for defense, showing moderate adaptability to threats but limited combat readiness overall.
  5. (5)
    Predictable Behavior: Episodes are characterized by high consistency in resource collection and basic tool crafting actions, with low variance across episodes.
  6. (6)
    Inconsistent Health and Exploration Management: While the agent effectively manages drink levels, it shows less consistency in food and health management and sacrifices exploration efficiency for achievement unlocking.
Crafter – Random
  1. (1)
    Ineffective Crafting and Resource Management: The agent frequently attempts crafting without the necessary resources or understanding of prerequisites, leading to repeated failures and minimal progress in achieving complex objectives.
  2. (2)
    Poor Survival Strategy: The agent consistently demonstrates ineffective survival behavior, including health depletion and poor management of food and drink levels, which hampers its ability to sustain itself in the game environment.
  3. (3)
    Limited Achievement Progression: While the agent reliably unlocks basic achievements like wake_up and collect_sapling, it struggles to achieve more complex milestones that require crafting, resource management, or combat engagement.
  4. (4)
    Repetitive and Ineffective Actions: Episodes are marked by high frequencies of ‘noop‘ actions and repetitive failed attempts at crafting, reflecting a lack of strategic adaptation and learning from past failures.
  5. (5)
    Lack of Combat Engagement: The agent shows minimal engagement with combat mechanics and fails to defend effectively against threats such as zombies and skeletons.
  6. (6)
    Predictable Behavior: Across episodes, the agent exhibits consistent, ineffective patterns of action, suggesting significant limitations in its decision-making processes and adaptability.
[…]
Insights from agents’ SySLLM summaries in the MiniGrid environments.
Figure 4. Insights from agents’ SySLLM summaries in the MiniGrid environments.
Together, these examples demonstrate that abstractive textual summaries can function as faithful and interpretable accounts of reinforcement learning policies.
7. Expert Evaluation of Summaries
We complement the quantitative performance analysis with an expert-based evaluation of SySLLM summaries. The goal is to assess how well the generated summaries capture the behavioral motifs observed by human experts and to quantify correctness while identifying potential hallucinations.
Evaluation Protocol. We recruited six graduate students with research experience in the training and evaluation of RL agents. The experts were divided into two groups: Experts 1–3 annotated the MiniGrid-Unlock (goal-directed and short-sighted) agents, while Experts 4–6 annotated the remaining MiniGrid agents. In addition, Experts 1, 2, and 6 annotated the Crafter agents. Each expert was shown a 120-second video per agent, depicting representative trajectories. Based on these trajectories, the experts were instructed to produce textual summaries using the same SySLLM prompting guidelines (see Appendix G). This alignment ensures comparability between expert- and model-generated summaries.
Scoring Framework. To compare SySLLM summaries $S_{\text{LLM}}$ with expert summaries $\{S_{\text{exp}}^{j}\}_{j=1}^{m}$, we decompose both into sets of atomic key points, denoted $\mathcal{K}_{\text{LLM}}$ and $\mathcal{K}_{\text{exp}}^{j}$, respectively. Semantic equivalence is assessed at the level of atomic propositions, where each key point expresses a single behavioral claim. Following standard practice in evaluating human-interpretable explanations (hoffman2018metrics), propositions are manually extracted and matched without embedding-based or heuristic similarity measures. Each SySLLM key point is labeled as a full match, partial match, or unsupported relative to the expert set. This procedure yields a transparent and reproducible matching process, with substantial inter-annotator agreement (AC1 = 0.72). The full annotation protocol is provided in the appendix.
Each pairwise comparison is scored as:
$\text{match}(k_{\text{LLM}},k_{\text{exp}})=\begin{cases}1&\text{if semantically equivalent},\\ 0.5&\text{if partially overlapping},\\ 0&\text{otherwise}.\end{cases}$
Recall. Recall measures the extent to which the model summary covers expert-identified key points:
$\text{Recall}=\frac{1}{m}\sum_{j=1}^{m}\frac{\sum_{k\in\mathcal{K}_{\text{exp}}^{j}}\max_{k^{\prime}\in\mathcal{K}_{\text{LLM}}}\text{match}(k,k^{\prime})}{\lvert\mathcal{K}_{\text{exp}}^{j}\rvert}$.
Precision. Precision measures the correctness of SySLLM key points relative to expert judgments. For each expert $j$, annotators are shown the set difference $\mathcal{K}_{\text{LLM}}\setminus\mathcal{K}_{\text{exp}}^{j}$ and asked to label each key point as Matched, Partially Matched, or Not Matched. Precision is defined as:
$\text{Precision}=\frac{\sum_{k\in\mathcal{K}_{\text{LLM}}}\max_{j}\text{match}(k,\mathcal{K}_{\text{exp}}^{j})}{\lvert\mathcal{K}_{\text{LLM}}\rvert}$.
This formulation captures both coverage (recall) and correctness (precision), while enabling explicit identification of hallucinated content through unmatched model key points.
Results. Table 3 reports per-agent recall and precision. Recall scores range from 0.687 (Unlock goal-directed) to 0.914 (Crossing), with a mean of 0.840, demonstrating substantial overlap with expert-identified points. Precision scores range from 0.769 (Dynamic Obstacles) to 0.864 (Short-Sighted), with a mean of 0.839, indicating minimal hallucination. For example, in the Crossing environment, the point “The agent frequently checks for walls in its path and adjacent tiles” was rejected by all experts, illustrating a rare hallucination.
Inter-Annotator Agreement. To ensure the reliability of expert annotations, we calculated both raw agreement (percentage of identical match scores across experts) and Gwet’s AC1 coefficient, which is robust to class imbalance in categorical judgments. Across all agents, the mean raw agreement was 70%, while the mean AC1 reached 0.72, indicating substantial inter-rater reliability. This ensures that the observed recall/precision metrics are not artifacts of inconsistent annotations.
Taken together, these results show that SySLLM summaries exhibit both high coverage and correctness relative to expert annotations, with recall and precision consistently above 0.8.
Agent Expert Recall Precision Mean
Unlock Goal-Directed E1 0.500 0.864 R = 0.687, P = 0.864
E2 0.643 0.864
E3 0.917 0.864
Unlock Short-Sighted E1 0.800 0.846 R = 0.878, P = 0.839
E2 0.833 0.807
E3 1.000 0.923
Dynamic Obstacles E4 0.583 0.692 R = 0.739, P = 0.769
E5 0.833 0.692
E6 0.800 0.923
Lava Gap E4 0.667 0.769 R = 0.794, P = 0.811
E5 0.786 0.846
E6 0.929 0.818
Red-Blue Doors E4 0.857 0.767 R = 0.871, P = 0.834
E5 0.857 0.867
E6 0.900 0.867
Crossing E4 0.750 0.731 R = 0.914, P = 0.795
E5 0.917 0.808
E6 1.000 0.846
Crafter Resource-Collector E7 0.938 0.893 R = 0.931, P = 0.871
E8 0.938 0.857
E9 0.917 0.864
Crafter Random E7 0.929 0.917 R = 0.902, P = 0.929
E8 1.000 0.958
E9 0.778 0.857
Table 3. Per-expert recall and precision scores for SySLLM summaries, with aggregated per-agent means. $R$ = Recall, $P$ = Precision. Overall averages across all agents: $R=0.840$, $P=0.839$.
8. User Study
@@
We conducted a controlled user study to evaluate the usefulness of SySLLM summaries compared to HIGHLIGHTS-DIV (HIGHLIGHTS) amir2018highlights, a standard demonstration-based benchmark in XRL which selects a set of high importance and diverse execution trajectories. The study assessed both subjective preferences and objective task performance. We focus on three qualitatively distinct agents from the MiniGrid Unlock environment (goal-directed, short-sighted, and random), ensuring the diversity of policies from structured strategies to noisy behaviors. For completeness, we provide a short description of HIGHLIGHTS in Appendix A.
Experimental Design. The study used a mixed design with two tasks. Task 1 (Preferences) followed a within-subject setup: each participant evaluated both modalities (SySLLM and HIGHLIGHTS), with order counterbalanced to mitigate ordering effects. Task 2 (Identification) followed a between-subject setup: Participants viewed only one modality, aligned with their Task 1 order. This produced four experimental conditions that varied by summary modality and agent type, as shown in Table 4. For HIGHLIGHTS, highlight videos were generated using 300 traces, a context length of 5, and 20 highlights, while SySLLM summaries were generated from 50 captioned episodes (see Section 5).
@@ 4 HIGHLIGHTS $\rightarrow$ SySLLM Short-sighted Agent HIGHLIGHTS
Table 4. Experimental conditions.
Procedure. The participants first completed a tutorial on MiniGrid Unlock rules, followed by a comprehension quiz. In Task 1, they watched a 120 seconds video of an agent’s behavior and then rated a summary (SySLLM or HIGHLIGHTS) on eight explanation quality metrics (7-point Likert), adapted from Hoffman et al. hoffman2018metrics. After evaluating both modalities, they provided direct preference judgments: which summary better reflected the agent’s policy and by what margin. In Task 2, participants were shown a summary (textual or visual) and asked to match it to one of three short (20s) videos: the correct agent plus two distractors. Each participant completed three trials ($Q1$: goal-directed, $Q2$: random, $Q3$: short-sighted). For each, they indicated their choice, rated confidence, and provided a justification.
Participants. We recruited 200 participants from Prolific (native English speakers from the US, UK, Canada, and Australia). Compensation was £3.75 base plus a £1 bonus for correct completion of Task 2. After exclusions for failed attention checks and implausibly short completion times (below 300 seconds), 192 participants remained (94 female, $M_{age}=36.4$, $SD=12.1$).
Results. In Task 1, SySLLM consistently outperformed HIGHLIGHTS across all metrics (Fig. 5). Paired $t$-tests confirmed the difference as highly significant ($T=13.99$, $p<10^{-33}$). Direct preference questions reinforced this: 75.5% of participants favored SySLLM, and the comparative Likert rating averaged $M=5.97$, $SD=1.44$ (neutral baseline = 4). Qualitative feedback highlighted that SySLLM explained why agents acted as they did, while HIGHLIGHTS required subjective inference. For example: “There are instances in the video where the agent seems to turn random corners. The summary explains why.” In Task 2, correctness rates for both modalities exceeded random-guess baselines (Fig. 6). Chi-Square tests found no significant differences between SySLLM and HIGHLIGHTS across Q1–Q3. However, confidence scores revealed a significant effect in Q3, where participants in the SySLLM condition reported higher confidence ($t=3.42$, $p=0.0008$).
Overall, participants rated SySLLM summaries as significantly clearer and more informative than the highlight videos. Although both modalities supported correct agent identification, textual summaries provided stronger interpretive cues, particularly reflected in higher confidence for certain agents. These results suggest that abstractive, language-based policy summaries enhance subjective interpretability while maintaining competitive performance in behavior recognition.
[…]
Participant scores for Task 1.
@@
Figure 5. Participant ratings for Task 1 on a 1–7 Likert scale. SySLLM ratings are significantly higher than HIGHLIGHTS ratings across all metrics.
[…]
Correctness rate in Task 2.
Figure 6. Correctness rate in Task 2. Error bars indicate 95% confidence intervals. Both SySLLM and HIGHLIGHTS outperform the random guess baseline. The differences in correctness rate between SySLLM and HIGHLIGHTS is not statistically significant.
9. Discussion and Future Work
We introduced SySLLM, a framework for abstractive textual explanation of reinforcement learning policies. By converting state and action trajectories into structured language and leveraging large language models to generate policy summaries, SySLLM abstracts low-level decision traces into high-level behavioral patterns. Empirical results show close alignment with expert interpretations and a clear user preference over visual demonstrations, supporting language-based abstraction for interpreting complex policies.
Several limitations remain. SySLLM currently relies on domain-specific captioning functions to translate trajectories into text. While feasible in simulated environments, extending to high-dimensional or partially observed domains will require perceptual grounding via visual-language pipelines or pretrained vision-language models capable of zero-shot scene understanding.
Future work extends beyond static summarization. One direction is interactive policy querying, where users ask natural language questions about agent behavior, enabling counterfactual and rationale-based explanations. Another is comparative summarization, contrasting policies across training stages, reward functions, or architectures. Finally, integrating textual and visual modalities and advancing toward open-domain policy summarization may enable benchmark-agnostic systems that characterize agent behavior at scale.
{acks}
Funded by the European Union (ERC, Convey, 101078158) and the Israel Science Foundation (ISF) under Grant No. 3079/24. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.